# The Role of Task Complexity in Emergent Abilities of Small Language Models

### v1

@@ 4 Results
Appendix A Additional plots
A.1 Modulo 26
Figure 8: PCA of embeddings: We choose all models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 9: PCA of embeddings: We choose all models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
A.2 Modulo 10
Figure 11: PCA of embeddings: We choose all models which reached over $90$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure MAX, or all mixtures involving a given operation, e.g. Mixed SUM. Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do. Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
Figure 12: PCA of embeddings: We choose all models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 13: PCA of embeddings: We choose all models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
@@ A.2 Modulo 10
Figure 17: PCA of embeddings with the same sample size (67): We choose the smallest 67 models which reached over $95$% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. We observe that the PCs get slightly distorted with the smaller sample size compared to fig. 13, but the overall structure stays the same. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 20: Evolution of training loss, accuracy, and PCs of cosine similarities in the embedding layer. ListOps base 10. The top main figure shows the evolution of training loss (solid lines) and test accuracy (dashed lines) for models with an embedding dimension of 128 and 3 layers, trained either on SUM-only data (blue) or on mixed MAX+MED+SUM data (red). Curves represent the mean across three independent runs, with shaded regions indicating one standard deviation. All models were trained for 20000 iterations. The red and blue boxes beneath the main plot display PCs revealing that models trained on MAX+MED+SUM data progressively develop a structured representation of numerical concepts, accompanied by a steady decrease in loss. In contrast, models trained solely on SUM data exhibit no clear structure in the embedding space and show long plateaus in the loss curve.
Appendix B Triplet Experiments
To increase control in our experimental setup, we train models only on triplet inputs 21. We construct 1000 unique triplet samples, splitting them into 900 for training and 100 for testing.
Appendix C Related works
Scaling laws for language models characterize how validation loss scales with model size (number of parameters) hoffmann2022training, compute (FLOPs) muennighoff2024scaling, dataset size (number of examples) hestness2017deep, and information capacity (in bits) allen2024physics, offering a foundation for model design. However, this perspective overlooks critical aspects of learning. Our results show that the composition of the training data—specifically, the mix of tasks—can significantly alter what and how a model learns. Even models with identical sizes can follow different learning trajectories depending on the training setup, suggesting that task structure shapes the internal algorithms that emerge during training.
Joint training Joint training on compositional tasks allows models to acquire the foundational primitives necessary for solving complex operations, often leading to abrupt performance improvements once all constituent skills are learned lubana2024percolation; okawa2023compositional. Similarly, training on multiple arithmetic tasks has been shown to improve accuracy on individual operations, highlighting the benefits of shared representations for compositional generalization lee2024teaching.
Number representation Neural networks trained on modular addition tasks develop structured embedding spaces that reflect underlying arithmetic operations. Mechanistic analyses have shown that small models can exhibit ordered number patterns and implement distinct algorithmic strategies, depending on initialization and hyperparameters zhong2023clock. Some models converge to known solutions such as the Clock algorithm, while others discover novel procedures like the Pizza algorithm, illustrating the algorithmic diversity that can emerge from fixed training data. Periodic structures in the embeddings can be characterized via Fourier analysis, offering additional interpretability nanda2023progress. These behaviors have also been linked to grokking dynamics, where models abruptly generalize after extended training, accompanied by the emergence of structured embedding patterns power2022grokking; liu2022towards.
In general language models performs poorly on symbolic mathematical tasks frieder2024mathematical; dziri2024faith; dave2024investigating such as the ListOps dataset nangia2018listops used in this study. Models often struggle with generalization, tending to memorize tasks rather than simulate the underlying algorithms. While mathematical tasks prove challenging for large language models to learn, they provide a controllable playground to test how models learn different tasks and to evaluate their accuracy quantitatively. Furthermore it allows us to tune the task difficulty by combining different operations.
Here, we investigate how language models acquire arithmetic skills by training small models on the ListOps dataset, which enables explicit evaluation through structured mathematical expressions. Adopting a bottom-up approach, we find that models learn to solve these tasks once the number of trainable parameters surpasses a critical threshold. This threshold shifts with the difficulty of the target operation (MAX = MIN < MED < SUM), indicating a dependency on task difficulty. Surprisingly, joint training on multiple operations often facilitates learning, outperforming models trained on individual operations (e.g. MAX+MED+SUM < SUM). This suggests that task diversity can ease optimization by promoting shared representations.
Appendix D Data and processing
D.1 Dataset Notation
[…]
D.2 Tokenization
We employ a character-based tokenization strategy for processing ListOps expressions. This approach offers several advantages:
@@ D.2 Tokenization
    Simplicity: Character-level tokenization eliminates the need for complex tokenization rules or a large vocabulary.
@@
    Generalizability: It allows the model to potentially generalize to unseen number combinations or deeper nesting levels.
Each character in the ListOps expression, including digits, operation symbols, and structural elements (parentheses, ’s’, ’e’), is treated as a separate token. This granular representation enables the model to learn the syntactic structure of the expressions alongside their semantic content.
D.3 Chain of Thought Implementation
We find that directly solving nested ListOps in one step can be quite challenging for transformer model (Fig. 22) Even with a maximum of three nesting levels with three operands (inputs) we find that GPT models with over 10 million parameters still fail to learn the task. To enhance model performance, particularly on more complex operations like sum modulo 10, we introduced a chain of thought (CoT) approach in our training data. This method involves providing step-by-step solutions that resolve the deepest nesting level at each step. For example:
s(%12(%34))>(%127)>0=0e
In this CoT representation:
  • The initial expression is s(%12(%34))
  • The first step resolves the innermost operation: (%34) becomes ‘7‘
  • The intermediate result is shown: s(%12(7))>(%127)
  • The process continues until the final result is reached: s(%12(7))>(%127)>0=0e
This CoT approach serves multiple purposes: 1. It guides the model through the problem-solving process, mimicking human-like reasoning. 2. It provides more granular supervision, potentially aiding in learning complex operations. 3. It allows us to study how models learn to break down and solve nested problems. Our experiments show that this CoT method significantly improves model performance, particularly for the challenging sum modulo 10 operation (Fig. 22).
Appendix E Observations from the norm of Attention and Feedforward outputs
@@
    Attention ratio: $r_{attn}=\frac{\|SA(LN_{1}(x))\|}{\|x\|}$
  2. 2.
    Feedforward ratio: $r_{ffwd}=\frac{\|FFN(LN_{2}(x_{1}))\|}{\|x_{1}\|}$
Attention Sublayer
  • Kolmogorov-Smirnov test: statistic = 0.1592, p-value < 0.0001
  • Jensen-Shannon divergence: 0.1591
  • Wasserstein distance: 0.0696
  • Effect size (Cohen’s d): 0.1761
  • 95% CI for mean difference: (0.0466, 0.0860)
Feedforward Sublayer
The feedforward sublayer exhibited more pronounced differences:
  • Kolmogorov-Smirnov test: statistic = 0.2461, p-value < 0.0001
  • Jensen-Shannon divergence: 0.1617
  • Wasserstein distance: 0.2830
  • Effect size (Cohen’s d): -0.3379
  • 95% CI for mean difference: (-0.3042, -0.2226)
Interpretation
  1. 1.
  2. 2.
  3. 3.
Appendix F Ablation studies
Figure 24: Random sum table. Joint training helps in learning even randomized sum tables. The numbers in the sum table follow a uniform distribution. We found that a randomized sum table with a non-uniform distribution is also easy to learn. However, a randomized sum table with the same uniform distribution remains as hard to learn as the original sum table. Early stopping criteria were applied during training in all simulations.
Figure 25: Whittaker-Shannon number representation for cyclic SUM Modulo 10. Top row: Cyclic matrices generated using Whittaker–Shannon interpolation illustrate how digit embeddings can encode modular addition via matrix multiplication. Middle row (red box): Embeddings learned by a model trained on MAX+MED+SUM. Bottom row (blue box): Cyclic SUM representation constructed using Whittaker–Shannon numbers and digit ordering, showing a similar structure to the model’s embeddings—suggesting that the model discovers an emergent encoding of cyclic SUM.
Appendix G Complexity: Number of operands and nesting level.
In previous studies, task complexity has been characterized through various measures, including the number of bits required to memorize the task, which corresponds to the length of the expression dave2024investigating, the number of operands and nesting depth petruzzellis2024benchmarking, and the structure of the computational graph dziri2024faith, which captures the number of nestings. Here, we examine how the number of operands and nesting levels affect the learning ability of small language models. Focusing on the all3 task, which combines max, med, and sum operations, we manipulate complexity by varying the number of operands (arg = {3, 4, 5}) and nesting depth (depth = {3, 4, 5}). The length of equations, measured by the number of characters, depends on both the number of operands and the nesting level. We find that nesting has a greater impact: increasing the nesting level from 3 to 4 results in longer equations than increasing the number of operands from 3 to 5 at a fixed nesting level. We train the GPT model on the all3 dataset with all combinations of arg (number of operands) and depth (nesting levels), finding that the model’s performance correlates with the sum of operands and nesting levels (arg + depth). Notably, transition points tend to group together for configurations with the same sum (Fig. 26, 27b). While Fig. 27a demonstrates that the model requires more parameters to solve longer equations, it also indicates that arg + depth serves as a reliable predictor of the transition point.
(a) Average equation length vs. transition point.
(b) Transition points.
Appendix H Experiment settings.
  1. 1.
    Model: GPT model (karpathy2022nanogpt)
@@ Appendix F Ablation studies
Appendix J Shuffled SUM modulo 26
@@
Appendix K Attention and Feedforward outputs modulo 10
Figure 32: Attention layer and Feedforward layer ratio histogram. Each row shows the attention and feedforward layer ratio histogram for models trained on MAX, MED, SUM, MAX+MED,and MAX+MED+SUM. Each column shows the ratio histogram in different attention blocks. The model had 128 embedding and 3 layers and all models rechec 99% accuracy.

### v2

@@ Appendix A Related works
Appendix B Additional plots
B.1 Modulo 20
Figure 10: ADD+PROD embeddings and PCs. PCs colored based modulo 4.
Figure 11: ADD, PROD, ADD+PROD embeddings and PCs.
B.2 Modulo prime - PROD
Figure 12: PCs PROD modulo prime vs not prim.
Appendix C Permutation groups.
Figure 13: Permutation - block diagonal with two 3x3 block matrix Red indicates a cycle sub-block, while blue indicates a non-cycle block.
Figure 14: 6x6 block diagonal permutation operations(OP, Op_TOP, OP_Bottom) OP_Top only acts on the top block matrix and OP_Bottom only acts on the bottom block matrix, while OP is an operation that acts on both the top and bottm blocks.
Figure 15: Permutation data structure. Left: Operation table (OP) - block diagonal with two 3x3 block matrix. Middle: Operation table (OP_BOTTOM) - block diagonal with two 3x3 block matrix. Right: Operation table (OP_TOP) - block diagonal with two 3x3 block matrix
Figure 16: Permutation experiments: The correlation matrix and top 2 PCA embedding f=of permutaion matrices for high performing models in two settings: Left: trianing on the full $6\times 6$ matrix multiplication (OP). Right: trianing on a mixture of OP, and subtasks OP_BOTTOM (multiplying bottom $3\times 3$ block, repeating the top block of first matrix), and and OP_TOP (multiply top $3\times 3$ block, repeat bottom of first matrix). The mixed task model reaches $90\%$ accuracy at 1.1k parameters, whereas the single OP model requires 1.9k for the same accuracy level.
Appendix D Shuffled SUM
The Shuffled SUM table was constructed using the following code:
Listing 1: Generate upper triangle and diagonal matrices
# Generate upper triangle and diagonal matrices.
# To ensure commutativity i+j=j+i, we will transpose the
# the upper triangle.
upper_triangle_matrix = {(i, j): (i + j) % MOD
for i in range(MOD)
for j in range(MOD) if i < j}
# To ensure uniform distribution after shuffling the values,
# we must shuffle the diagonal separately.
# This is because all off-diag r.h.s. are repeated twice
# for commutativity, but not the diagonal entries.
diagonal_matrix = {(i, i): (2 * i) % MOD for i in range(MOD)}
# Function to shuffle values in a dictionary
def shuffle_dict_values(d):
keys = list(d.keys())
values = list(d.values())
random.shuffle(values)
return dict(zip(keys, values))
# Apply shuffling
shuffled_triangle = shuffle_dict_values(upper_triangle_matrix)
shuffled_diagonal = shuffle_dict_values(diagonal_matrix)
As a result, there is no consistent mapping between the original and shuffled values (e.g., the number 1 does not always map to 2), making it difficult for the model to learn a deterministic transformation. While the shuffled sum table remains commutative, we did not enforce associativity. A direct check also confirmed that it does not satisfy the associative property for most triplets.
Figure 17: Shuffle the summation table.
D.1 Shuffled SUM modulo 10
@@
D.2 Shuffled SUM modulo 26
@@
Appendix E Deep transformer vs. Recurrent Transformer
Figure 23: Recurrent Transformer vs. Deep transformer In the Rec-GPT block, there is only one transformer block that is applied recursively n times, so each output becomes the input for the next step. In contrast, in a traditional GPT block, there are n transformer blocks stacked sequentially on top of each other.
Appendix F Results using Deep Transformer based models
F.1 Experiment settings
@@
    Model: GPT model ((Karpathy, 2022))
@@
F.3 Data and processing
F.3.1 Dataset Notation
max(3,min(7,4,9))=4 $\quad\Rightarrow\quad$ Polish: (max,3,(min,7,4,9))=4
To disentangle any complexity arising from tokenization we further simplify these expression by representing the by symbols: ’++’ for max, ’-’ for min, ’//’ for median, and $\%$’ for sum modulo 10. For example:
(max,3,(min,7,4,9))=4 $\quad\Rightarrow\quad$ Our notation: s(+3(-749))=4e
F.3.2 Tokenization
We employ word-level tokenization where each semantically meaningful unit is a single token:
  • Operations: Each operation with its modulus is one token (e.g., add_20, max_10, prod_19).
  • Numbers: Each number is one token, regardless of digit count (e.g., 3, 13, 104 are each single tokens).
  • Symbols: Parentheses (, ), equals =, CoT separator >, and end token <eos> are each single tokens.
Note that commas are not needed as token boundaries are implicit. For modulo $n$, the vocabulary size is approximately $n$ + (number of operations) + 5 (for symbols). This tokenization ensures that each number is atomic, preventing the model from learning spurious digit-level patterns.
F.3.3 Chain of Thought Implementation
Directly solving nested ListOps in one step is challenging for transformers (Fig. 24). Even with maximum three nesting levels and three operands, GPT models with over 10 million parameters fail to learn. To address this, we use a chain-of-thought (CoT) approach that resolves the innermost operation at each step.
For example, with the functional notation:
add_10(1 2 add_10(3 4))>add_10(1 2 7)>0=0<eos>
The CoT proceeds as:
  • Initial expression: add_10(1 2 add_10(3 4))
  • Step 1: Resolve innermost add_10(3 4) $\to$ 7
  • Step 2: Resolve add_10(1 2 7) $\to$ 0
  • Final: 0=0<eos>
This CoT approach guides the model through step-by-step reasoning, provides granular supervision, and significantly improves performance on challenging operations (Fig. 24).
F.4 Modulo 26
Figure 25: PCA of embeddings: We choose all models which reached over 9595% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 26: PCA of embeddings: We choose all models which reached over 9595% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 28: PCA of embeddings: We choose all models which reached over 9090% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure MAX, or all mixtures involving a given operation, e.g. Mixed SUM. Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do. Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
F.5 Modulo 10
Figure 31: PCA of embeddings: We choose all models which reached over 9090% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure MAX, or all mixtures involving a given operation, e.g. Mixed SUM. Interestingly, pure SUM does not show a discernible structure in the embeddings, whereas all other cases do. Notably, Mixed SUM models exhibit a prominent odd-even separation in PC5.
Figure 32: PCA of embeddings: We choose all models which reached over 9595% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 33: PCA of embeddings: We choose all models which reached over 9595% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
Figure 35: PCA of embeddings with the same sample size (67): We choose the smallest 67 models which reached over 9595% test accuracy. Each row shows the average correlation matrix and top PCs for models trained on either a single operation, e.g. Pure+MAX, or all mixtures involving a given operation, e.g. Mixed+MAX. We observe that the PCs get slightly distorted with the smaller sample size compared to fig. 33, but the overall structure stays the same. Again, pure SUM does not show a discernible structure in the embeddings, whereas all cases do.
F.6 Triplet Experiments
To increase control in our experimental setup, we train models only on triplet inputs 38. We construct 1000 unique triplet samples, splitting them into 900 for training and 100 for testing.
Appendix G Observations from the norm of Attention and Feedforward outputs
@@
    Feedforward ratio: $r_{ffwd}=\frac{\|FFN(LN_{2}(x_{1}))\|}{\|x_{1}\|}$
Attention Sublayer
  • Kolmogorov-Smirnov test: statistic = 0.1592, p-value ¡ 0.0001
  • Jensen-Shannon divergence: 0.1591
  • Wasserstein distance: 0.0696
  • Effect size (Cohen’s d): 0.1761
  • 95% CI for mean difference: (0.0466, 0.0860)
Feedforward Sublayer
The feedforward sublayer exhibited more pronounced differences:
  • Kolmogorov-Smirnov test: statistic = 0.2461, p-value ¡ 0.0001
  • Jensen-Shannon divergence: 0.1617
  • Wasserstein distance: 0.2830
  • Effect size (Cohen’s d): -0.3379
  • 95% CI for mean difference: (-0.3042, -0.2226)
Interpretation
  1. 1.
  2. 2.
  3. 3.
Figure 40: Attention patterns and layer dynamics in SUM vs MAX+MED+SUM (all3) models. Each panel shows (CoT) solution to a SUM modulo 10 problem, where ’¿’ indicates solution steps. The first row shows the input sequence, with curved lines representing attention weights from Layer 2 in a 3-layer network. Black lines highlight attention patterns for a specific digit (shown in orange). Below are shown various layer metrics including the ratio of self-attention to feedforward norms (sa/ffwd), self-attention output norms (sa_out), feedforward output norms (ffwd_out), and ratios of layer outputs to inputs (sa_out/x, ffwd/x). Top: Model trained only on SUM operations shows attention primarily focused on parentheses and structural elements. Bottom: Model trained on MAX+MED+SUM (all3) shows attention strongly connecting to digits being combined in each CoT step, suggesting direct involvement in numerical computation. These distinct patterns suggest fundamentally different algorithms learned by each model.
Figure 41: Attention layer and Feedforward layer ratio histogram. Each row shows the attention and feedforward layer ratio histogram for models trained on MAX, MED, SUM, MAX+MED,and MAX+MED+SUM. Each column shows the ratio histogram in different attention blocks. The model had 128 embedding and 3 layers and all models reaches 99% accuracy.
@@
Appendix H Ablation Studies
H.1 Number of Attention Heads
[…]
H.2 Deep vs. Recurrent Architecture
Our primary architecture uses a recurrent transformer: a single transformer block applied iteratively, feeding its output back as input for the next iteration. We compared this to standard deep transformers with 2–4 stacked layers and found qualitatively similar results for the task difficulty hierarchy and joint training benefits (see Section F). The recurrent architecture was chosen for two reasons: (1) it converges faster during training, and (2) it produces more structured, interpretable embeddings that facilitate our PCA analysis. The key finding—that joint training with compatible tasks reduces model size requirements—holds for both architectures.
Appendix I Additional Scaling Results
Figure 44 shows the full scaling results across all task combinations tested, complementing the selected comparisons in the main text.
Figure 44: Full scaling results across moduli. Transition points (minimum model size to learn) as a function of modulo $n$ for various task combinations. Prime moduli are shown in white text. Key observations: (1) PROD-prime is consistently hard; (2) ADD+PROD helps ADD for non-prime moduli; (3) MAX+PROD helps PROD even for primes; (4) ADD+MAX+PROD at primes can be harder than individual tasks.
Appendix J Additional Curriculum Experiments
Figure 45 shows the full curriculum training results including loss, accuracy, and mix ratio curves comparing baseline ADD training with the ADD+PROD $\to$ ADD curriculum.
Figure 45: Full curriculum comparison: ADD+PROD $\to$ ADD. Left: Validation loss over training. Middle: Accuracy over training—baseline ADD (blue) stays at $\sim$15%, curriculum (orange) reaches near-perfect accuracy. Right: Curriculum mix ratio showing transition from ADD+PROD to pure ADD.
In addition, we tested transfer from MAX+MED to ADD.
MAX+MED $\to$ ADD transfer.
We trained a small model ($n_{\text{embed}}=48$, below the pure ADD transition) on MAX+MED until proficiency, then gradually transitioned to pure ADD (never showing mixed MAX+MED+ADD expressions). Key findings:
  • The model learns ADD immediately upon introduction, achieving high accuracy.
  • The model forgets MAX and MED as they are phased out, but retains the structured embedding.
[…]
This confirms that the curriculum learning effect is robust across different task combinations: structured representations learned from easy tasks enable efficient learning of hard tasks.
Figure 46: Transfer from MAX+MED to ADD. A model too small to learn pure ADD ($n_{\text{embed}}=48$) successfully learns ADD after pretraining on MAX+MED. Blue: curriculum (MAX+MED $\to$ ADD) reaches high accuracy. Red: baseline (pure ADD) fails to learn.
Appendix K Complexity: Number of operands and nesting level.
In previous studies, task complexity has been characterized through various measures, including the number of bits required to memorize the task, which corresponds to the length of the expression (Dave et al., 2024), the number of operands and nesting depth (Petruzzellis et al., 2024), and the structure of the computational graph (Dziri et al., 2023), which captures the number of nestings. Here, we examine how the number of operands and nesting levels affect the learning ability of small language models. Focusing on the all3 task, which combines max, med, and sum operations, we manipulate complexity by varying the number of operands (arg = {3, 4, 5}) and nesting depth (depth = {3, 4, 5}). The length of equations, measured by the number of characters, depends on both the number of operands and the nesting level. We find that nesting has a greater impact: increasing the nesting level from 3 to 4 results in longer equations than increasing the number of operands from 3 to 5 at a fixed nesting level. We train the GPT model on the all3 dataset with all combinations of arg (number of operands) and depth (nesting levels), finding that the model’s performance correlates with the sum of operands and nesting levels (arg + depth). Notably, transition points tend to group together for configurations with the same sum (Fig. 47, 48b). While Fig. 48a demonstrates that the model requires more parameters to solve longer equations, it also indicates that arg + depth serves as a reliable predictor of the transition point.
(a) Average equation length vs. transition point.
(b) Transition points.
Appendix L Clustering in the embedding space
Figure 49: ADD, PROD and ADD+PRD in MOD 42 PC embeddings
Figure 50: ADD, PROD and ADD+PRD in MOD 43 PC embeddings
Figure 51: ADD, PROD and ADD+PRD in MOD 45 PC embeddings
Figure 52: Clustering in the PC space. We measure how well modulo-2, 3, 4, 5, and 6 divisor clusters separate in the PC space for the ADD, PROD, and ADD+PROD models on MOD 42, 43, and 45 (for runs with accuracy $80\%$). Cluster quality is computed from the mean position and standard deviation of each modulo group, and the pairwise distances between them. For MOD 42, strong clustering appears for mod-2, 3, 6, and sometimes 4—precisely the divisors of 42. Strong PROD signals almost always produce strong ADD+PROD signals, whereas strong ADD signals without PROD support do not transfer. For MOD 43, no model shows strong clustering; both ADD and PROD struggle and do not help each other. For MOD 45, strong clustering occurs for mod-3 and mod-5, and occasionally mod-6 (via divisor 3). We also observe that the mixed model only exhibits strong clustering when PROD does as well, indicating that the PROD embedding drives the shared structure and helps ADD.
Figure 53: Clustering in the PC space by averaging over the first 5 PCs. We measure how well modulo-2, 3, 4, 5, and 6 divisor clusters separate in the PC space for the ADD, PROD, and ADD+PROD models on MOD 42, 43, and 45 (for runs with accuracy $80\%$). Cluster quality is computed from the mean position and standard deviation of each modulo group, and the pairwise distances between them, averaged over the first five PCs. For MOD 42, strong clustering appears for mod-2, 3, and 4 (could be an artifact of 2) — precisely the divisors of 42. For MOD 43, we find no strong clustering compared to MOD 42 and MOD 45. For MOD 45, strong clustering occurs for mod-3 and mod-5, as expected from the divisors of 45. We also observe that the mixed model exhibits strong clustering only when PROD does as well, indicating that the PROD embedding drives the shared structure and supports ADD.